OpenStudy (anonymous):

Determine where Q(t)=1/2t-ln(1+t^2) is not changing.

OpenStudy (campbell_st):

differentiate and let the derivative = 0 and solve for t \[dy/dx = 1/2 - 2t/(1+t^2)\] so (1 + t^2) = -4t or t^2 +4t + 1=0 Use the general quadratic formula to solve \[t = (-4 \pm \sqrt{4^2 - 4 \times1 \times 1})/(2\times1)\] Its not changing when \[t = -2 \pm \sqrt{3}\]

OpenStudy (anonymous):

ooh, i didn't expect that the process will be like this. .

OpenStudy (anonymous):

Wow thanks, now let me get to it until i can understand.